# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v2

@@ Abstract
Recent advances in multimodal foundation models have set new standards in few-shot anomaly detection. This paper explores whether high-quality visual features alone are sufficient to rival existing state-of-the-art vision-language models. We affirm this by adapting DINOv2 for one-shot and few-shot anomaly detection, with a focus on industrial applications. We show that this approach does not only rival existing techniques but can even outmatch them in many settings. Our proposed vision-only approach, AnomalyDINO, is based on patch similarities and enables both image-level anomaly prediction and pixel-level anomaly segmentation. The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning. Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$). The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, e.g., in industrial contexts.
@@ 1 Introduction
Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD). We collect the nominal patch representations from the (potentially augmented) reference sample(s) in the memory bank $\mathcal{M}$. At test time, we select the relevant patch representation via masking (if applicable). The distances of those to the nominal representations in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score $s(\mathbf{x}_{\mathrm{test}})$ using the aggregation statistic $q$. For both, masking and feature extraction, we utilize DINOv2. Further examples for other categories are depicted on the right (and in Figures 4 and 5 in Appendix A).
@@
Our approach, termed AnomalyDINO, is memory bank-based and leverages DINOv2 [29] as a backbone (but the approach could be adapted to other architectures with strong patch-level feature extraction). We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model). At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by [7] or [20]. Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD [2] and outperforms all but one competing method on VisA [48].
@@ 2 Related Work
Multimodal foundation models have emerged as powerful tools for a wide range of tasks, see e.g., [3, 22, 6, 16, 30, 24, 28]. Most relevant to visual AD are multimodal approaches based on CLIP [30] or recent LLMs [28], but also vision-only approaches like DINO [5, 29]. CLIP [30] learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations. The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs. This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts. DINO [5, 29] leverages a self-supervised student-teacher framework based on vision transformers [12]. It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks. DINOv2 [29] combines ideas from DINO with patch-level reconstruction techniques [46] and scales to larger architectures and datasets. The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training. GroundingDINO [24], builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
@@
However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples. Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD. The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets [17], Wide ResNets [44], and Vision Transformers (ViT) [11]. These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set. The comparison with the memory bank is typically conducted using a nearest neighbor approach [1, 8, 10, 33, 41, 23]. Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank. Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification. The underlying idea consists of two steps. First, these approaches define sets of prompts describing nominal samples and anomalies. Second, the corresponding textual embeddings are compared against the image embeddings [18, 7, 20, 47]. Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous. However, these methods either require significant prompt engineering (e.g., [7] use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings). Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots. These methods require more elaborate prompting and techniques for interpreting textual outputs [42]. Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
@@ 3 Matching Patch Representations for Visual Anomaly Detection
This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features. The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data. We first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$. Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$. A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
Inspired by PatchCore [33], our method simplifies the memory bank framework by selecting DINOv2 as the backbone, thereby reducing the complexity of the feature engineering stage—specifically in deciding which representations or layers to use and how to aggregate them. We also tailor the memory bank concept to the few-shot learning regime by introducing an improved aggregation statistic, and a new preprocessing pipeline that incorporates zero-shot masking and augmentations. The proposed method is described in detail in the following subsections.
@@ 3.1 Anomaly Detection via Patch (Dis-) Similarities
Let us briefly review the idea of the memory bank approach. We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}\in\mathcal{X}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{X}$ denotes some space of images and $\mathcal{F}$ the feature space. Note that $n$ depends on the image resolution and the patch size. Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in[k]\bigr{\}}$ (with shorthand $[k]:=\{1,\dots,k\}$), we collect the nominal patch features and store them in a memory bank
@@
To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with $\mathcal{M}$. To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
@@
Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$. The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile [26] and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal. Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated). However, one could also replace $q$ with a different statistic to cater to special cases. If anomalies are expected to cover larger parts of the image, for example, a high percentile of the patch distances might be a suitable choice. If, on the contrary, anomalies may occur only locally such that only very few patches/or a single patch might be affected, the score $q$ should be sensitive to the highest patch distances. Frequently, the maximum pixel-wise anomaly score is considered. Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation. Following [33], we utilize bilinear upsampling and Gaussian smoothing ($\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects. Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 4 and 5).
@@ 3.2 Enriching the Memory Bank & Filtering Relevant Patches
In the few-shot anomaly detection setting, the primary challenge is to effectively capture the concept of normality from the limited set of nominal samples. A useful strategy involves applying simple augmentations, like rotations (following the insights in [41]), to enhance the variability of nominal patch features. This is essential because the variability at test time is likely to be significantly greater than in the reference data, $X_{\mathrm{ref}}$.
To avoid irrelevant areas of the test image from leading to falsely high anomaly scores, we propose masking the object of interest. This approach mitigates the risk of false positives, particularly in the few-shot regime where the limited reference samples may not adequately capture the natural variations in the background, as exemplified in Figure 8, Appendix C.
It is important to note that the appropriate preprocessing technique—whether to apply masking and/or augmentations—should depend on the specific characteristics of the object of interest. For a more detailed discussion on the challenges and considerations in designing an effective preprocessing pipeline, see Section C.1.
@@
Masking, i.e., delineating the primary object(s) in an image from its background, can help to reduce false-positive predictions. To minimize the overhead of the proposed pipeline, we utilize DINOv2 also for masking. This is achieved by thresholding the first PCA component of the patch features [29]. We observe empirically that this sometimes produces erroneous masks. Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches. To address this issue, we check if the PCA-based mask accurately captures the object in the first reference sample and apply the mask accordingly, which gives rise to the ‘masking test’ in Figure 2. This test is performed only once per object as the procedure yields very consistent outputs. In addition, we utilize dilation and morphological closing to eliminate small holes and gaps within the predicted masks. See Figures 12 and 13 for examples of this masking procedure, Table 8 for the outcomes of the masking test per object, and Figure 8 for a visualization of the benefits of masking in the presence of background noise (all in Appendix C). In general, we do not mask textures (e.g., ‘Wood’ or ‘Tile’ in MVTec-AD).
@@ 4 Experiments
We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE [8], PaDiM [10], PatchCore [33], WinCLIP+ [18], and APRIL-GAN [7]. It is important to note that ACR [21] and MuSc [23] consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO. We adapt AnomalyDINO to this setting in Appendix D. Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method. In Tables 2 and 3, results reported by [18] are indicated by , the result of the WinCLIP re-implementation by [20] (where J. Jeong is also coauthor) by . All other results are taken from the original publication. For ADP [20] we report the (usually slightly better) variant ADP (with access to the class label). For GraphCore no standard deviations are reported [41].
@@ Results
We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively.
@@
Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in{1,2,4,8,16}$ for every reported metric, outperforming approaches that require additional training data sets (such as ADP and APRIL-GAN). The method also demonstrates superior anomaly localization, with results showing that while detection performance remains comparable across different resolutions (448 vs. 672), a higher resolution enhances localization performance. Furthermore, we observe clear improvements as the number of samples increases.

### v3

@@ Abstract
Recent advances in multimodal foundation models have set new standards in few-shot anomaly detection. This paper explores whether high-quality visual features alone are sufficient to rival existing state-of-the-art vision-language models. We affirm this by adapting DINOv2 for one-shot and few-shot anomaly detection, with a focus on industrial applications. We show that this approach does not only rival existing techniques but can even outmatch them in many settings. Our proposed vision-only approach, AnomalyDINO, follows the well-established patch-level deep nearest neighbor paradigm, and enables both image-level anomaly prediction and pixel-level anomaly segmentation. The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning. Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$). The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, e.g., in industrial contexts.
@@ 1 Introduction
footnotetext: This paper was accepted at the Winter Conference on Applications of Computer Vision (WACV), 2025. The final version is available here (Open Access) and on IEEE Xplore.
@@
Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD). We collect the nominal patch representations from the (potentially augmented) reference sample(s) in the memory bank $\mathcal{M}$. At test time, we select the relevant patch representation via masking (if applicable). The distances of those to the nominal representations in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score $s(\mathbf{x}_{\mathrm{test}})$ using the aggregation statistic $q$. For both, masking and feature extraction, we utilize DINOv2 ($f$). Further examples for other categories are depicted on the right (and in Figures 4, 5, 6 and 7 in Appendix A).
@@
Our approach, termed AnomalyDINO, follows the well-established AD framework of patch-level deep nearest neighbor [46, 34], and leverages DINOv2 [30] as a backbone. We carefully design a suitable preprocessing pipeline for the few-shot scenario, using the zero-shot segmentation abilities of DINOv2 (which alleviates the additional overhead of another segmentation model). At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank $\mathcal{M}$.
Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as [7] or [21]. Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD [2] and outperforms all but one competing method on VisA [51].
@@ 2 Related Work
Multimodal foundation models have emerged as powerful tools for a wide range of tasks, see e.g., [3, 23, 6, 17, 31, 25, 29]. Most relevant to visual AD are multimodal approaches based on CLIP [31] or recent LLMs [29], but also vision-only approaches like DINO [5, 30]. CLIP [31] learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations. The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs. This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts. DINO [5, 30] leverages a self-supervised student-teacher framework based on vision transformers [12]. It employs a multi-view strategy to train Vision Transformers (ViT) [11] to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks. DINOv2 [30] combines ideas from DINO with patch-level reconstruction techniques [49] and scales to larger architectures and datasets. The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training. GroundingDINO [25], builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
@@
A well-established approach is that of deep nearest neighbor AD, where test instances are scored according to the distance to their nearest neighbors (in feature space) from the memory bank $\mathcal{M}$, containing features of nominal instances. This approach dates back to (at least) 2002 [13], but got later popularized first on image-level [1] and then by the seminal works on patch-level: Patch support vector data description (SVDD) [46] trains a patch-based feature extractor (with a mixture of DeepSVDD [37] and a self-supervised loss), and PatchCore [34] utilizes pre-trained classification models on Imagenet for patch-level feature extraction. Various further approaches build on this simple, but effective AD approach [8, 43, 24]. Besides kNN matching, other ‘classical’ AD methods like the Mahalanobis distance are employed [10]. Evidently, the performance of these approaches crucially depends on the feature extractor $f$. Classical choices are ResNets [18], Wide ResNets [47], and Vision Transformers [11], mostly pre-trained on a supervised task. Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification. The underlying idea consists of two steps. First, these approaches define sets of prompts describing nominal samples and anomalies. Second, the corresponding textual embeddings are compared against the image embeddings [19, 7, 21, 50]. Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous. However, these methods either require significant prompt engineering (e.g., [7] use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings). Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots. These methods require more elaborate prompting and techniques for interpreting textual outputs [44]. Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
@@ 3 Matching Patch Representations for Visual Anomaly Detection
This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features. We build upon the well-established deep nearest neighbor approaches [13, 46, 8, 34, 43, 1, 24], i.e., we first gather relevant patch representations of nominal patches in a memory bank $\mathcal{M}$. Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$. A suitable aggregation of the patch-based distances gives anomaly scores on image-level.
Our work differs from previous deep nearest neighbor approaches [34, 46] by tailoring the memory bank concept to the few-shot regime, specifically utilizing the strong features from DINOv2: We design a pipeline that incorporates zero-shot masking and augmentations, and we propose a more robust aggregation statistic. Importantly, we identify DINOv2 as the ideal backbone for our scenario due to its strong patch-level features and masking ability. In addition, this simplifies the deep nearest neighbor framework by reducing the complexity of the feature engineering stage (e.g., by making it unnecessary to think of which representations/layers to use and how to aggregate them), and increasing the flexibility w.r.t. input resolution (that can be chosen to be any multiple of $14$). The proposed method is described in detail in the following subsections.
@@ 3.1 Anomaly Detection via Patch (Dis-) Similarities
Let us briefly review the idea of patch-level deep nearest neighbor AD. We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}\in\mathcal{X}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{X}$ denotes some space of images and $\mathcal{F}$ the feature space. Note that $n$ depends on the image resolution and the patch size (in the case of DINOv2, a patch is $14\times 14$ pixels). Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\{\mathbf{x}^{(i)}\mid\,i\in[k]\}$ (with shorthand $[k]:=\{1,\dots,k\}$), we collect the nominal patch features and store them in a memory bank
@@
To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with $\mathcal{M}$. To do so, we leverage a nearest neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
@@
Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in the set $\mathcal{D}$. The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile [27] and turns out to be suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal. Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated). However, one could also replace $q$ with a different statistic to cater to special cases. If anomalies are expected to cover larger parts of the image, for example, a high percentile of the patch distances might be a suitable choice. If, on the contrary, anomalies may occur only locally such that only very few patches/or a single patch might be affected, the score $q$ should be sensitive to the highest patch distances. Frequently, the maximum pixel-wise anomaly score is considered. Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation.
Following [34], we utilize bilinear upsampling and Gaussian smoothing ($\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects. Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 4, 5, 6 and 7).
@@ 3.2 Enriching the Memory Bank & Filtering Relevant Patches
In the few-shot anomaly detection setting, the primary challenge is to effectively capture the concept of normality from the limited set of nominal samples. A useful strategy involves applying simple augmentations, like rotations (following the insights in [43]), to enhance the variability of nominal patch features. This is essential because the variability at test time is likely to be significantly greater than in the reference data, $X_{\mathrm{ref}}$. In contrast, full-shot methods aim to reduce the size of $\mathcal{M}$ to reduce inference time, e.g., [34].
To avoid irrelevant areas of the test image from leading to falsely high anomaly scores, we propose masking the object of interest. This approach mitigates the risk of false positives, particularly in the few-shot regime where the limited reference samples may not adequately capture the natural variations in the background, as exemplified in Figure 10, Appendix C. It is important to note that the appropriate preprocessing technique—whether to apply masking and/or augmentations—should depend on the specific characteristics of the object(s) of interest. For a more detailed discussion on the challenges and considerations in designing an effective preprocessing pipeline, see Section C.1.
@@
Masking, i.e., delineating the primary object(s) in an image from its background, can help to reduce false-positive predictions, thereby improving the robustness in the low-data regime. To minimize the overhead of the proposed pipeline, we utilize DINOv2 also for masking. This is achieved by thresholding the first PCA component of the patch features [30]. We observe empirically that this sometimes produces erroneous masks. Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches. To address this issue, we check if the PCA-based mask accurately captures the object in the first reference sample and apply the mask accordingly, which gives rise to the ‘masking test’ in Figure 2. This test is performed only once per object as the procedure yields very consistent outputs. In addition, we utilize dilation and morphological closing to eliminate small holes and gaps within the predicted masks. See Figures 14 and 15 for examples of this masking procedure, Table 8 for the outcomes of the masking test per object, and Figure 10 for a visualization of the benefits of masking in the presence of background noise (all in Appendix C). In general, we do not mask textures (e.g., ‘Wood’ or ‘Tile’ in MVTec-AD).
@@ 4 Experiments
We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE [8], PaDiM [10], PatchCore [34], WinCLIP+ [19], and APRIL-GAN [7]. It is important to note that ACR [22] and MuSc [24] consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO. We adapt AnomalyDINO to this setting in Appendix D. Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method. In Tables 2 and 3, results reported from [19] are indicated by , the result of the WinCLIP re-implementation [21] (where J. Jeong is also coauthor) by . All other results are taken from the original publication. For ADP [21] we report the (usually slightly better) variant ADP (with access to the class label). For GraphCore no standard deviations are reported [43].
@@ Results
We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively. Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in\{1,2,4,8,16\}$ for every reported metric, outperforming approaches that require additional training data sets (such as ADP and APRIL-GAN). The method also demonstrates superior anomaly localization, with results showing that while detection performance remains comparable across different resolutions (448 vs. 672), a higher resolution enhances localization performance. Furthermore, we observe clear improvements as the number of samples increases.